A hamburger and a glass of soda cost $2.10. Two hamburgers and 3 sodas cost $5.15. Find the cost of each.
Found 2 solutions by ewatrrr, Gersid:
Answer by ewatrrr(24785)


Hi
Let x and y represent the cost of a hamburger anda soda respectively
Question states***
x + y = $2.10
2x + 3y = $5.15
-2x - 2y = -$4.20
2x + 3y = $5.15
y = $0.95,the cost of a soda. A hamburger costs $1.15 (2.10-.95)
CHECKING our Answer***
2*1.15 + 3*.95 = $2.30 + $2.85 = $5.15

Answer by Gersid(33)

Let H = number of hamburgers and S = number of sodas.
1) H+S = $2.10 "A hamburger and a glass of soda cost $2.10."
2) 2H+3S = $5.15 "Two hamburgers and 3 sodas cost $5.15."
Solve equation 1) for H and substitute this into equation 2).
1a) H = $2.10-S
2a) 2($2.10-S)+3S = $5.15 Simplify and solve for S.
2b) $4.20-2S+3S = $5.15 Combine the S terms and then subtract $4.20 from both sides.
2c) S = $0.95 One glass of soda costs $0.95
From 1a):
H = $2.10-S
H = $2.10-$0.95
H = $1.15 One hamburger costs $1.15